# Reinforcement Learning for Fine-tuning Text-to-speech Diffusion Models

### v1

@@ 5.1 Experimental Design
Figure 3: (fine-tuning effectiveness) The relative effectiveness of seven RL algorithms. (Left) shows the change of UTMOS during training with 550 reward samples (reward samples = episodes * batch size) of each RL approach while (right) shows the change of UTMOS during the initial 200 training episodes of the diffusion model loss-guided policy optimization (DLPO), which illustrates an increase of UTMOS from 3.41 to 3.59. $\textrm{DPOK}_{\textrm{tail}}$ and $\textrm{KLinR}_{\textrm{tail}}$ are only trained with samples of step $t=996,997,998,999$.
@@
We use the unofficial implementation of WaveGrad2 as our baseline model, pre-trained on the LJSpeech dataset [15] which consists of 13,100 short audio clips of a female speaker and the corresponding texts, totaling approximately 24 hours. For the reward model, we use UTMOS [29], trained on datasets from the VoiceMOS Challenge 2022 [14], including 4,974 audio clips from male and female speakers with MOS ratings. For automatic evaluation of the fine-tuned models, we use another speech quality and naturalness assessment model (NISQA) [21], trained on the NISQA Corpus including more than 14,000 speech samples produced by male and female speakers along with samples’ MOS ratings. (This use of a separate MOS network is intended to guard against overfitting the reward model.) We also conduct a human experiment to have people evaluate the speech quality of the generated audios from fine-tuned WaveGrad2.
@@
We follow Fan et al. [8] and use online RL training for fine-tuning WaveGrad2 to evaluate the performance of RWR, DDPO, DLPO, KLinR, $\textrm{KLinR}_{\textrm{tail}}$, DPOK, and $\textrm{DPOK}_{\textrm{tail}}$. In each episode (amount of batch processed) during the training, we sample a new trajectory based on the current model distribution $\pi_{\theta}$ and calculate the rewards of the new trajectory. Online RL training is claimed to be better at maximizing the reward than the supervised approach which only uses the supervised dataset based on the pre-trained distribution. Due to significant computational and memory demands of online RL training for fine-tuning WaveGrad2, we train WaveGrad2 with 8 A100-SXM-80GB GPU for 5.5 hours and 550 samples are fed to WaveGrad2.
During online RL fine-tuning, the model is updated using new samples from the previously trained model. In every training episode, the UTMOS score is computed using the final state $x_{0}$ of the trajectory generated given sampled text $c$ from the training dataset, where $c\sim p(c)$. We plot the UTMOS score obtained in each training episode in Figure 3 to illustrate the fine-tuning progress of Wavegrad2. In Figure 3 (left) shows the UTMOS score obtained during 550 training episodes for each RL approach. For Wavegrad2 baseline model trained with RWR, DDPO, DLPO, KLinR, and DPOK, we follow Wavegrad2 by randomly sampling step $t$ from 1000 diffusion/denoising steps during training, while for Wavegrad2 baseline model trained with $\textrm{DPOK}_{\textrm{tail}}$ and $\textrm{KLinR}_{\textrm{tail}}$, we only sample step $t$ from steps 996, 997, 998, 999, as we find that the KL divergence of current distribution and the distribution of previously trained model differs greatly in the last few steps. As presented in Figure 3 (left) DLPO increases UTMOS in the first 200 training episodes, shown in Figure 3 (right), and then maintains the UTMOS score around 3.4 for the rest of the samples during the training, while other methods have UTMOS score drop to below 1.5 in the first 200 training episode and then maintain the UTMOS score. Moreover, the $x_{0}$ generated by methods other than DLPO gradually becomes acoustically noisy as the number of seen samples grows. Due to randomness in sampled text, different models receive different initial samples, resulting in varied UTMOS at the beginning, as shown in Figure 3.
Figure 4: (a) shows the mean UTMOS and NISQA scores for audio based on 200 unseen texts generated by models fine-tuned using seven RL algorithms, the error bar shows the standard deviation of each result. (b) shows the proportion of raters who prefer the audios generated from the DLPO fine-tuned model or baseline model and the proportion of raters who think audios generated by DLPO fine-tuned model and baseline model are about the same (Tie).
5.2 Algorithm Comparisons
@@
We save the top three checkpoints for each model and use them to generate audios for 200 unseen texts. We then use UTMOS and NISQA to predict MOS score of these generated audios. The results are shown in Figure 4(a).
We see that only audios generated by DLPO fine-tuned Wavegrad2 gain UTMOS and NISQA scores higher than the audios generated by the baseline pretrained WaveGrad2 model. Two two-sample t-tests confirmed significant differences ($p<10^{-20}$) between the baseline and DLPO for both UTMOS and NISQA scores. For models trained with random sampling steps, RWR outperforms DDPO, KLinR and DPOK in fine-tuning TTS diffusion models, though its UTMOS and NISQA scores are below the baseline. Nonetheless, it is observed that audios generated through methods other than DLPO exhibit significant fricative noise, indicating lower speech quality. We observe that the models trained with tail steps sampling, both $\textrm{DPOK}_{\textrm{tail}}$ steps and $\textrm{KLinR}_{\textrm{tail}}$ steps, generate audio with higher UTMOS and NISQA scores compared to the model trained with DPOK and KLinR using random sampling. The detailed results of UTMOS and NISQA scores for each checkpoints and links to sample audios are shown in Appendix C.
@@
We further conduct an experiment recruiting human participants to evaluate the speech quality of the audios generated by DLPO. We randomly select 20 audios pair from the generated audios (among audios of 200 unseen texts). Each pair includes one generated audio from DLPO and one from the baseline pretrained WaveGrad2 model, both audios based on the same text. 11 listeners are recruited for the experiment. They are asked to assess which audio is better regarding to speech naturalness and quality. We show the results in Figure 4. In 67% of comparisons, audios generated from DLPO fine-tuned Wavegrad2 is rated as better than audios generated by the baseline Wavegrad2 model, while 14% of comparisons have audios generated by the baseline Wavegrad2 model rated as better. 19% of comparisons are rated as about the same (Tie). The experiment question is shown in Appendix E.
@@ 6 Discussion
Our research indicates that fine-tuning TTS diffusion models using DLPO results in better outcomes compared to other reinforcement learning methods. This is likely due to DLPO aligning with the training procedure of TTS diffusion models by incorporating the original diffusion model loss as a penalty in the reward function to effectively prevent model deviation in fine-tuned TTS models. As demonstrated by Ouyang et al. [24], RL fine-tuning can be improved by constraining the model to avoid arbitrary deviations from the original model. However, DDPO, DPOK and KLinR do not improve Wavegrad2 in our experiments despite the latter two incorporating KL divergence terms intended to provide this constraint. The formulas used to calculate loss in these methods incorporate a KL upper-bound estimate [8] which we must further approximate in training due to memory constraints, by sampling both trajectories and then inference steps $t$. These approximations are very unstable in practice due to the large number of steps in each inference trajectory (in Appendix D).
Moreover, as demonstrated in our experiments, models trained with KLinR_tail and DPOK_tail generate audios with higher UTMOS and NISQA scores compared to KLinR and DPOK. This suggests that certain inference steps are more crucial for the performance of the fine-tuned model, but approximating KL by sampling steps may miss these. Previous studies did not encounter this issue because the text-to-image diffusion model they used had fewer inference steps and DDPO, DPOK, and KLinR for images could calculate KL based on all inference steps. It would likely be possible to trade time for performance by using more samples to obtain better KL bounds which would allow stable fine-tuning with these methods. However, we show that the original loss term (DLPO) is the only one which remains stable enough for fine-tuning with single samples. Future work should explore more efficient methods for approximating KL in models with long inference chains.
Appendix A Algorithm for fine-tuning Wavegrad2 with DPOK
@@
     Sample step $t\sim\mathcal{U}\{0,T\}$, compute $\log p_{\theta}(x_{t-1}|x_{t},c)$ and $\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}$
     Update gradient using Equation 10 and update $\theta$
@@
Appendix B Algorithm for fine-tuning Wavegrad2 with KLinR
@@
     Sample step $t\sim\mathcal{U}\{0,T\}$, compute $\log p_{\theta}(x_{t-1}|x_{t},c)$ and $\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}$
     Update gradient using subsection 4.4 and update $\theta$
@@
Appendix C Fine-tuned Results: UTMOS and NISQA scores
RL Algorithms $\alpha$ $\beta$ UTMOS NISQA
baseline - - 2.90 3.74
RWR 1 - 2.18 3.00
RWR 5 - 1.67 2.95
RWR 0.5 - 2.50 3.42
RWR 0.1 - 1.46 2.84
DDPO 1 - 1.30 2.30
DDPO 5 - 1.20 1.09
DLPO 1 1 3.52 3.70
DLPO 5 1 3.47 3.92
DLPO 10 1 3.31 3.63
DLPO 1 0.8 3.71 3.96
DLPO 1 0.6 3.53 3.70
DLPO 1 0.5 3.61 3.83
KLinR 1 1 1.45 2.33
KLinR 1 10 1.43 0.82
KLinR 10 1 1.37 1.46
KLinR 1 0.1 1.51 1.47
KLinR 1 0.01 1.42 1.29
$\textrm{KLinR}_{\textrm{tail}}$ 1 1 1.51 2.84
$\textrm{KLinR}_{\textrm{tail}}$ 1 0.1 1.51 2.68
$\textrm{KLinR}_{\textrm{tail}}$ 1 0.01 2.60 3.62
$\textrm{KLinR}_{\textrm{tail}}$ 1 0.001 1.89 2.78
DPOK 1 1 1.41 2.83
DPOK 1 10 1.34 2.55
DPOK 1 100 1.29 1.17
DPOK 10 1 1.37 1.46
$\textrm{DPOK}_{\textrm{tail}}$ 1 1 1.51 2.84
$\textrm{DPOK}_{\textrm{tail}}$ 1 0.1 1.51 2.68
$\textrm{DPOK}_{\textrm{tail}}$ 1 0.01 1.70 3.62
$\textrm{DPOK}_{\textrm{tail}}$ 1 0.001 1.98 2.78
Table 1: Mean NISQA mos and naturalness scores for generated audios
Sample audios are presented in this website https://anonymous.4open.science/r/audios-C7F7/
Appendix D Experiment of KL divergence stability
We run two experiments to evaluate the stability of KL divergence estimation. We first use two fixed checkpoints (baseine checkpoint vs. checkpoint finetuned with DPOK) to calculate the KL divergence of samples at selected inference step between two fixed checkpoints, we refer to as Monte Carlo KL. Then we calculate the KL divergence of the entire audio between two fixed checkpoints, which we refer to as KL upperbound. The results are shown in the following figure. We can see that both Monte Carlo KL and KL upperbound estimation are very unstable in practice, and the KL upperbound has more outliers may due to the large number of steps in estimation.
Figure 5: shows KL divergence of samples at selected inference steps from fixed policy trained with DPOK and fixed baseline policy, labeled as MonteCarlo_KL; exact KL upperbound of samples from fixed policy trained with DPOK and fixed baseline policy, labeled as KL_upperbound.
Appendix E Human Experiment Example

### v2

@@ 5.1 Experimental Design
1 Reward UTMOS During Training
2 Evaluation NISQA MOS During Training
@@
Figure 1: (fine-tuning effectiveness) The relative effectiveness of five RL algorithms. (Top) shows the the change of reward UTMOS during initial 160 training episodes (training samples = episodes * batch size) of each RL approach while (bottom) shows the change of evaluation NISQA MOS during the same initial 160 training episodes. Left figures shows the different methods’ training performance and right figures shows that DLPO increases UTMOS from 3.0 to 3.68 and increases NISQA from 3.85 to 4.12.
Dataset
WaveGrad 2 is pre-trained on the LJSpeech dataset [15] which consists of 13,100 short audio clips of a female speaker and the corresponding texts, totaling approximately 24 hours. We use WaveGrad 2’s training set and validation set to finetune Wavegrad 2 (12388 samples for training, 512 samples for validation. 200 unseen samples are used as test set for evaluation.
Evaluation Metrics
@@
For automatic evaluation of the fine-tuned models, we use another pretrained speech quality and naturalness assessment model (NISQA) [24], trained on the NISQA Corpus including more than 14,000 speech samples produced by male and female speakers along with samples’ MOS ratings. (This use of a separate MOS network is intended to guard against overfitting the reward model.) We also conduct a human experiment to have people evaluate the speech quality of the generated audios from fine-tuned WaveGrad2. To evaluate the intelligibility of the synthesized audio, we transcribe the speech with a pre-trained ASR model, Whisper [32], and compute the word error rate (WER) between the transcribed text and original transcript.
RL Fine-tuning
@@
We follow Fan et al. [7] and use online RL training for fine-tuning WaveGrad2 to evaluate the performance of RWR, DDPO, DLPO, KLinR, DPOK. In each episode (amount of batch processed) during the training, we sample a new trajectory based on the current model distribution $\pi_{\theta}$ and calculate the rewards of the new trajectory. Online RL training is claimed to be better at maximizing the reward than the supervised approach which only uses the supervised dataset based on the pre-trained distribution. We train WaveGrad2 with 8 A100-SXM-80GB GPU for 5.5 hours. We set the batch size as 64 and the denoising steps as 10.
During online RL fine-tuning, the model is updated using new samples from the previously trained model. In every training episode, the UTMOS score is computed using the final state $x_{0}$ of the trajectory generated given sampled text $c$ from the training dataset, where $c\sim p(c)$, meanwhile the evaluation MOS score of $x_{0}$ is also calculated using NISQA.
We plot both UTMOS and NISQA MOS scores obtained in each training episode in Figure 1 to illustrate the fine-tuning progress of Wavegrad2. (Top) shows the UTMOS scores over 160 training episodes for each RL method. DLPO, DPOK, and KLinR initially increase UTMOS greatly during the first 40 episodes. DLPO then gradually rises above 3.6 and remains stable for the rest of the training, while DPOK stays around 3.5. KLinR reaches a score of 3.5 by episode 120 but starts to decline afterward. In contrast, both DDPO and RWR show a steady decrease in UTMOS from the start. By episode 100, DDPO’s UTMOS falls below 1.5, with only a slight increase afterward, while RWR’s UTMOS continues to drop, reaching below 1.5 by the end of the 160 episodes.
(Bottom) shows the evaluation NISQA MOS over 160 training episodes for each RL method. DLPO increases NISQA greatly from 3.85 to 4.12 during the first 100 episodes, then maintain above 4.05. KLinR also increases NISQA to above 4 but starts to decrease after 100 episodes. DPOK has NISQA decrease from 4.0 to 3.65 during the initial 60 episodes then gradually rises above 4.0. Both RWR and DDPO has steady decrease in NISQA from the start. Moreover, the $x_{0}$ generated by DDPO and RWR gradually becomes acoustically noisy as the number of seen samples grows. Due to randomness in sampled text, different models receive different initial samples, resulting in varied UTMOS and NISQA at the beginning.
Is diffusion loss alone effective?
We conduct an experiment that finetunes the baseline WaveGrad2 with only the diffusion loss as reward, which we name OnlyDL. The loss function is $-\log p_{\theta}(x_{t-1}|x_{t},c)*(-\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2})$. We train this model for 5.5 hours which is the same training time as DLPO and we plot the total loss, diffusion loss, and change of UTMOS during training in tensorboard, shown in Appendix E. Both total loss and diffusion loss has clearly plateaued and shows minimal fluctuations during training. Moreover, the UTMOS maintains around 3 during the entire training, and DLPO has UTMOS increase from around 3 to above 3.65 during training, which is reported in Figure 3 in our paper. This also indicates that DLPO can effectively increase base model’s UTMOS during training.
One vs. ten sampled steps in diffusion loss guidance
To further assess the influence of denoising steps, we conducted an experiment where we fine-tuned the baseline WaveGrad2 model using DLPO with only a single denoising step, in contrast to the previous experiment that used 10 denoising steps. In this experiment, we randomly select one denoising step $t\sim\mathcal{U}\{0,999\}$. Then we compute the diffusion model loss $\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2}$ and log probability $\log p_{\theta}(x_{t-1}|x_{t},c)$ based on $x_{t}$, update the gradient following Equation 10 and update $\theta$. The results are shown in Table 2.
5.2 Experiment Results
@@
We save the top three checkpoints for each model during training and use them to generate audios for 200 unseen texts. We then use UTMOS and NISQA to predict MOS score of these generated audios and the real human speech audios, labeled as ground truth. The results are shown in Table 1 and Figure 2(a). Sample audios are presented on https://demopagea.github.io/DLPO_demo/.
[…]
Ground Truth - - 4.20 4.37 -
Base Model - - 2.90 3.74 1.5
OnlyDL - - 3.16 3.45 1.4
RWR 1 - 2.18 3.00 8.9
DLPO 1 1 3.65 4.02 1.2
DDPO 1 1 2.69 2.96 2.1
KLinR 1 1 3.02 3.73 1.3
DPOK 1 1 3.18 3.76 1.1
Table 1: Mean UTMOS, NISQA MOS, and word error rate for generated audios. Ground Truth is the audio of real human speech.
RL Algorithms Denoising Steps UTMOS $\uparrow$ NISQA $\uparrow$ WER $\downarrow$
DLPO 1 3.71 3.96 1.7
DLPO 10 3.65 4.02 1.2
Table 2: NISQA, UTMOS, and WER for DLPO with different denoising steps.
Figure 2: (a) shows the mean UTMOS and NISQA scores for generated audio based on 200 unseen texts, the error bar shows the standard deviation of each result. (b) shows the proportion of raters who prefer the audios generated from the DLPO fine-tuned model or baseline model and the proportion of raters who think audios generated by DLPO fine-tuned model and baseline model are about the same (Tie).
We observe that the audios generated by the DLPO fine-tuned WaveGrad2 achieve the highest UTMOS score of 3.65 and the highest NISQA score of 4.02, both significantly better than those produced by the baseline pretrained WaveGrad2 model. Additionally, the WER for DLPO-generated audios is 1.2, lower than that of the baseline model. Two two-sample t-tests confirmed significant differences ($p<10^{-20}$) between the baseline and DLPO for both UTMOS and NISQA scores.
Both DPOK and KLinR also show effectiveness in improving base model. Both of them improve UTMOS, and DPOK also improves NISQA in this experiment; however, both DDPO and RWR fail to improve the base model and their generated audios are noisy.
When comparing DLPO with OnlyDL, DLPO outperforms OnlyDL in UTMOS, NISQA, and WER. This suggests that while diffusion gradients can assist in model fine-tuning, our RL method, DLPO, is much more effective in improving the base model performance, bringing it closer to ground truth quality. Additionally, as shown in Table 2, DLPO with varying denoising steps achieves similar UTMOS scores, whereas DLPO with 10 denoising steps exhibits improved performance in both NISQA and WER.
@@
We further conduct an experiment recruiting human participants to evaluate the speech quality of the audios generated by the previous version of DLPO. We randomly select 20 audios pair from the generated audios (among audios of 200 unseen texts). Each pair includes one generated audio from DLPO and one from the baseline pretrained WaveGrad2 model, both audios based on the same text. 11 listeners are recruited for the experiment. They are asked to assess which audio is better regarding to speech naturalness and quality. We show the results in Figure 2. In 67% of comparisons, audios generated from DLPO fine-tuned Wavegrad2 is rated as better than audios generated by the baseline Wavegrad2 model, while 14% of comparisons have audios generated by the baseline Wavegrad2 model rated as better. 19% of comparisons are rated as about the same (Tie). The experiment question is shown in Appendix F.
@@ 6 Discussion
In our experiments, we found that RWR and DDPO do not enhance TTS models as they do for text-to-image models, largely because they fail to effectively control the magnitude of deviations from the original model. Only methods that incorporate diffusion model gradients as a penalty are able to prevent such deviations. Both DPOK and KLinR, which apply diffusion model gradients as a regularized KL, succeed in reducing model deviation while improving the sound quality and naturalness of the generated speech. However, DLPO surpasses DPOK and KLinR, likely because it aligns more closely with the training process of TTS diffusion models by directly integrating the original diffusion model loss as a penalty in the reward function, thereby preventing model deviation in fine-tuned TTS models more effectively.
Additionally, the experiment comparing DLPO, which uses both human feedback and diffusion model gradients as rewards, with OnlyDL, which relies solely on diffusion model gradients to fine-tune the base diffusion model, reveals important insights. The results show that OnlyDL offers minimal improvement to the base diffusion model, while DLPO leads to significant enhancements. This suggests that diffusion model gradients alone may not be sufficient for substantial improvement in speech quality, perhaps because they fail to capture the complexities necessary for improving speech quality in TTS models. Human feedback, on the other hand, introduces an additional layer of guidance, steering the model toward more desirable outputs and effectively complementing the diffusion gradients. The success of DLPO demonstrates that integrating both human feedback and diffusion model loss as reward enables a more robust and efficient fine-tuning process, significantly enhancing the speech quality and naturalness of the generated output.
Comparing the performance of the WaveGrad2 model fine-tuned with a single denoising step to that of previous experiments using 10 denoising steps reveals that using more denoising steps improves speech quality and reduces word error rates. This enhancement is likely due to the fact that sampling additional denoising steps decreases variance and increases the chances of capturing more important denoising effects. The results suggest that employing multiple denoising steps enables a more thorough refinement of the generated audio, leading to clearer and more natural speech outputs. These findings underscore the significance of selecting appropriate denoising steps during the fine-tuning process, as they can greatly influence the model’s overall performance.
Appendix A DLPO implementation details
Figure 3: Detail steps for fine-tuning text-to-speech diffusion models with online RL learning
Pretrained Reward Model: we use UTMOS [36], a pretrained speech quality and naturalness assessment model to predict reward. The training dataset of UTMOS is provided by VoiceMOS Challenge 2022, which are collected by a large-scale MOS listening test from 288 human raters. UTMOS is trained to predict the human evaluation of the speech quality and naturalness (MOS).
Finetune WaveGrad2 with online learning RL: The implementation of online Reinforcement Learning (RL) involves several key components and steps that enable an agent to learn and adapt dynamically as it interacts with its environment. At its core, online RL requires a mechanism for real-time data collection, where the agent continuously observes the state of the environment, takes actions, and receives feedback in the form of rewards. In DLPO implementation, WaveGrad2 is updated based on the rewards and state transitions observed from the most recent updated WaveGrad2.
As shown in 3. More details for DLPO implementation, the most recent updated WaveGrad2 is WaveGrad2 $\theta$, we implement inference step of WaveGrad2 $\theta$ to generate Audio $x_{0}$ and the corresponding denoising trajectory of Audio $x_{0}$. Reward is calculated by using pretrained UTMOS to predict a MOS score for Audio $x_{0}$. Then we use the reward and denosing trajectory to update WaveGrad2 $\theta$ following DLPO algorithm.
Appendix B WaveGrad2 model
WaveGrad2 models the conditional distribution $p_{\theta}(y_{0}|x)$ where $y_{0}$ is the clean waveform and x is the conditioning features corresponding to $y_{0}$, such as linguistic features derived from the corresponding text:
$p_{\theta}(y_{0}|x):=\int p_{\theta}(y_{0:N}|x)\mathrm{d}y_{1:N}$ (12)
where $y_{1},...,y_{N}$ is a series of latent variables, each of which are of the same dimension as the data $y_{0}$, and N is the number of latent variables (iterations). The generative distribution $p_{\theta}(y_{0:N}|x)$ is called the denoising process (or reverse process), and is defined through the Markov chain, which is shown in Figure 4:
$p_{\theta}(y_{0:N}|x):=p(y_{N}))\prod_{n=1}^{N}p_{\theta}(y_{n-1}|y_{n},x)$ (13)
where each iteration is modelled as a Gaussian transition:
$p_{\theta}(y_{n-1}|y_{n},x):=\mathcal{N}\left(y_{n-1};\mu_{\theta}(y_{n},n,x),\textstyle\sum\nolimits_{\theta}(y_{n},n,x)\right)$ (14)
starting from Gaussian white noise $p(y_{N})=\mathcal{N}(y_{N};0,I)$. The approximate posterior $q(y_{1:N}|y_{0})$ is called the diffusion process (or forward process), and is defined through the Markov chain:
$q(y_{1:N}|y_{0}):=\prod_{n=1}^{N}q(y_{n}|y_{n-1})$ (15)
where each iteration adds Gaussian noise:
$q(y_{n}|y_{n-1}):=\mathcal{N}\left(y_{n};\sqrt{(1-\beta_{n})}y_{n-1},\beta_{n}I\right)$ (16)
under some (fixed constant) noise schedule $\beta_{1},...,\beta_{N}$. During training, we can optimize for the variational lower-bound on the log-likelihood (upper-bound on the negative log-likelihood):
$-\mathrm{log}p_{\theta}(y_{0}|x)\leq\mathbb{E}_{q}\left[-\log\frac{p_{\theta}(y_{0}|x}{q(y_{1:N}|y_{0})}\right]=\mathbb{E}_{q}\left[-\log p_{(}y_{N})-\sum_{n=1}^{N}\log\frac{p_{\theta}(y_{n-1}|y_{n},x)}{q(y_{n}|y_{n-1})}\right]$ (17)
According to Equation 17, a straightforward approach would be to parameterize a neural network to model the mean $\mu_{\theta}$ and variance $\Sigma_{\theta}$ of the Gaussian distribution described in Equation 7, allowing for direct optimization of the KL-divergence using Monte Carlo estimates. However, Ho et al. (2020) found it more effective to set $\Sigma_{\theta}$ as a constant following the $\beta_{n}$ schedule and reparameterize the neural network to model $\epsilon_{\theta}$, predicting the noise $\epsilon\sim\mathcal{N}(0,I)$ instead of $\mu_{\theta}$. With this reparameterization, the loss function can be expressed as:
Figure 4: WaveGrad2 follows a Markov process where forward diffusion $q(y_{n+1}|y_{n},x)$ iteratively adds Gaussian noise to the signal starting from the waveform $y_{0}$. $q(y_{t+1}|y_{0})$ is the noise distribution used for training. The inference denoising process progressively removes noise, starting from Gaussian noise $x_{T}$. This figure is adapted from Chen et al. [4], Ho et al. [11].
Appendix C Algorithm for fine-tuning Wavegrad2 with DPOK
@@
     Sample $x_{t}$ given $x_{0}$, $t\sim[0,T]$, compute $\log p_{\theta}(x_{t-1}|x_{t},c)$ and $\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}$
     Update gradient using subsection 4.4 and update $\theta$
@@
Appendix D Algorithm for fine-tuning Wavegrad2 with KLinR
@@
     Sample $x_{t}$ given $x_{0}$, $t\sim[0,T]$, compute $\log p_{\theta}(x_{t-1}|x_{t},c)$ and $\log\frac{p_{\theta}(x_{t-1}|x_{t},c)}{p_{\mathrm{pre}}(x_{t-1}|x_{t},c)}$
     Update gradient using Equation 9 and update $\theta$
@@
Appendix E Tensorboard plot diffusion loss only model
In this figure, diffusion model loss is labeled as l1loss (left), the loss function is $-\log p_{\theta}(x_{t-1}|x_{t},c)*(-\|\tilde{\epsilon}(x_{t},t)-\epsilon_{\theta}(x_{t},c,t)\|_{2})$ (middle). UTMOS is labelled as mosscore (right).
Appendix F Human Experiment Example